# Fragmentation in Collapsar Disks: Migration, Growth, and Emission

### v1

@@ 6 Conclusions
    Toomre instability ($Q<1$) is a generic feature of collapsar disks with high accretion rates ($\dot{M}\gtrsim 1M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$). For some combinations of $\{M,\alpha\}$, Toomre instability can emerge down to accretion rates of $\dot{M}\approx 0.2M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$. While these accretion rates are extreme, averaged over the lifetime of a long GRB, we emphasize that they only need to be sustained for short periods of time (fragmentation of a gravitationally unstable zone occurs in $\sim$ few ms, and inward migration of fragments in $\sim 100$ ms). Toomre instability generally arises in the outer, advective regions of collapsar disks, beyond a radius $R\sim 100r_{\rm S}$. However, for large central masses $M\gtrsim 30M_{\odot}$, a second zone of Toomre instability can emerge at very small radii as well.
@@
    Fragments migrate inwards rapidly, initially due to Type I torques. At small radii, Type I migration will eventually become unimportant, usually because the mass accumulation torque comes to dominate Type I migration. At the smallest radii ($r\lesssim\leavevmode\nobreak\ {\rm few}\leavevmode\nobreak\ \times R_{\rm g}$), GW losses will dominate the orbit evolution; at this stage, gaps sometimes open in the disk as well. The final outcome of fragment migration can be either tidal disruption or merger into the BH, depending on the fragment radius.
@@
    A major uncertainty in fragment evolution, beyond the scope of this paper to address, concerns the way in which gravitationally unstable fragments will undergo a contraction-collapse sequence to achieve hydrostatic equilibrium. As a limiting case, we consider a situation in which a ($M_{\rm f}\gtrsim 0.1M_{\odot}$) fragment is able to contract to the cold NS mass-radius relationship. This scenario sets a lower limit on the fragment tidal disruption radius and therefore an upper limit on (a) $\gamma$-ray variability from accretion of a disrupted fragment and (b) GW luminosity from fragment inspiral. Under this (optimistic) assumption, we find typical contributions to GRB variability, from the fragment tidal disruption, at the $\sim 10\%$ level, and peak GW strains of $h_{\rm GW}\sim 1.5\cdot 10^{-22}(d/100\leavevmode\nobreak\ {\rm Mpc})^{-1}$.
@@ Appendix A Clarification to the CB07 model
Where $\epsilon$ is the particle energy and $\mu$ is the chemical potential. For simplicity, we perform a change in variables to $\theta=k_{\rm b}T/m_{\rm e}c^{2}$, $\eta=\mu/k_{\rm b}T$ and $E$=$\epsilon/m_{\rm e}c^{2}$ that changes the FD distribution to:
@@ Appendix D Numerical method
$\begin{split}F^{+}-F^{-}=&u^{r}\left[\frac{U_{\rm i}H_{\rm i}-U_{\rm i-1}H_{\rm i-1}}{r_{\rm i}-r_{\rm i-1}}-\right.\vspace{2mm}\\ &\qquad-\frac{(U+P)}{\rho}\frac{\left(\rho_{\rm i}H_{\rm i}-\rho_{\rm i-1}H_{\rm i-1}\right)}{r_{\rm i}-r_{\rm i-1}}\biggr{]}\end{split}$ (45)
@@
Start by choosing: $\alpha$, $a$, $M$, $\dot{M}$, $r_{\rm out}$, an acceptable error ($N_{\rm er}$) and a numerical push ($N_{\rm p}$). Build a logarithmic grid of $N+1$ radial cells, where $r_{\rm i}$ ranges from $r_{\rm ISCO}=r_{\rm N}$ to $r_{\rm out}=r_{\rm 0}$ Solve for $\eta_{\rm e}$ and $\theta$ at $r_{\rm out}$ using the pressure equation, $Y_{\rm e}=0.5$ and $U/\rho=GM/r_{\rm out}$. Try solving for $\eta_{\rm e}(r_{\rm i}),$ $\theta(r_{\rm i})$, and $Y_{e}(r_{{\rm i}})$ at $r_{{\rm i}}$ with the $\eta_{\rm e}(r_{\rm i-1}),$ $\theta(r_{\rm i-1})$, and $Y_{\rm e}(r_{{\rm i-1}})$ from the previous $r_{{\rm i-1}}>r_{{\rm i}}$ using 3D root-finding, when solving for the pressure, energy and lepton conservation equations. Check whether root-finding worked by testing that for the lepton equation $\Big{|}\frac{RHS({\rm i})-LHS({\rm i})}{LHS({\rm i})}\Big{|}<N_{{\rm er}}$ Continue to the next step where ${\rm i=i+1}$. Set $Y_{\rm e}(r_{\rm i})=Y_{\rm e}(r_{\rm i-1})+N_{\rm p}$ and solve for $\eta_{\rm e}(r_{\rm i}),$ $\theta(r_{\rm i})$ using a 2D root-finding solving for the pressure and energy equations. Then use the $Y_{\rm e}(r_{\rm i})$ and the $\eta_{\rm e}(r_{\rm i}),$ $\theta(r_{\rm i})$ found in the 2D root-finding to solve for the next step. Where $\rm i=i+1$. Collect all the calculated $\eta_{\rm e}(r_{\rm i}),$ $\theta(r_{\rm i})$, and $Y_{e}(r_{{\rm i}})$ for all $r_{\rm i}$’s. Now we can calculate all of the disks properties, like temperature, density, scale height, pressure and more. While ${\rm i}\leq{\rm N}$       TrueWhen $\rm i>N$False
Figure 15: A flow chart of the numerical algorithm used to solve our model. RHS(i) and LHS(i) are the right and left-hand sides (respectively) of the lepton equation with values from the 3D root finding at step $r_{\rm i}$. $N_{\rm p}$ is a small numerical push set to $-10^{-15}$, which was needed because root finding had trouble at the outer regions where changes in $Y_{\rm e}$ are small. $N_{\rm er}=10^{-6}$.
@@ Appendix E Migration calculations
$f_{\rm\gamma}$ =$=\begin{cases}1&\tau_{{\rm\nu}}+\tau_{{\rm\widetilde{\nu}}}<1\\ \frac{\sqrt{x/2}+1/\gamma}{\sqrt{x/2}+1}&\tau_{{\rm\nu}}+\tau_{{\rm\widetilde{\nu}}}\geq 1\end{cases}$ (47c)
@@ F.1 GRB variability from tidal disruption of fragment
Figure 16: The ratio between the luminosity from the tidal disruption of the fragment to the steady state luminosity from BH accretion ($L_{\rm TD}/L_{\bullet}$). Left and right panels represent the same $\{M,\dot{M},\alpha\}$ parameter choices as in Fig. 2. White, black, gray and magenta squares are colored with the same color scheme as 11. Squares that: (1) have a number, represent the ratio between the luminosities as colored by the rainbow color system represented by the bar legend below the figure; (2) are filled in cyan represent that the tidal radius is smaller than the ISCO, meaning no TDE; (3) are filled in light grey, represent final fragment masses below the minimal stable mass for the cold EOS. The luminosity from the tidal disruption is roughly $\sim 10\%$ of the steady luminosity.
@@ F.5 Duration of $\dot{m}_{\rm f}>\dot{M}$
The fragment accretion rate $\dot{m}_{\rm f}$ commonly exceeds the BH accretion rate $\dot{M}$ for a short duration of time (as is visible in the main text, in Figs. 9 and 10). In this appendix we plot (in Fig. 28) two cross section of our parameter survey, in order to explore the duration of this period (which may be observable as extreme GRB variability, as a secondary jet from the accreting fragment may outshine the primary jet from the central BH). Fig. 28shows that this duration generally ranges between $0.5-11.2ms$.
@@ F.6 Extended parameter survey
Figure 29: An extended version of Fig. 6 in the main text, plotting the color-coded values for the largest initial fragment mass (represented as $\log_{10}(m_{\rm f,i}/M_{\odot})$) in the fragmenting $Q<1$ region, but now exploring the full 3 dimensional parameter space of $\{M,\dot{M},\alpha\}$. Each of the panels represents a different $\alpha$ value, ranging from $0.01-0.3$, while varying $\{M,\dot{M}\}$. Black, white, and grey squares represent breakdowns in model assumptions or solutions as in Fig. 4. We see that fragment initial masses span a wide range, from $4\times 10^{-4}\lesssim m_{\rm f}/M_{\odot}\lesssim 2\times 10^{-1}$. The initial masses tend to increase with declining accretion rate, or with increasing central mass or viscosity.
@@
Finally, in the extended survey of GW properties, we find little change in $f_{{\rm max}}$ (Fig. 32). Likewise, the peak strain $h_{{\rm max}}$ (Fig. 33) is qualitatively similar to the behavior seen in the main text, though the dependence on both $M$ and $\alpha$is now more clear: the peak GW strain grows modestly for increasing either $M$ or $\alpha$.

### v2

@@ 6 Conclusions
    Toomre instability ($Q<1$) is a generic feature of collapsar disks with high accretion rates ($\dot{M}\gtrsim 1M_{\odot}\penalty 10000\ {\rm s}^{-1}$). For some combinations of $\{M,\alpha\}$, Toomre instability can emerge down to accretion rates of $\dot{M}\approx 0.2M_{\odot}\penalty 10000\ {\rm s}^{-1}$. While these accretion rates are extreme, averaged over the lifetime of a long GRB, we emphasize that they only need to be sustained for short periods of time (fragmentation of a gravitationally unstable zone occurs in $\sim$ few ms, and inward migration of fragments in $\sim 100$ ms). Toomre instability generally arises in the outer, advective regions of collapsar disks, beyond a radius $R\sim 100r_{\rm S}$. However, for large central masses $M\gtrsim 30M_{\odot}$, a second zone of Toomre instability can emerge at very small radii as well.
@@
    Fragments migrate inwards rapidly, initially due to Type I torques. At small radii, Type I migration will eventually become unimportant, usually because the mass accumulation torque comes to dominate Type I migration. At the smallest radii ($r\lesssim\penalty 10000\ {\rm few}\penalty 10000\ \times R_{\rm g}$), GW losses will dominate the orbit evolution; at this stage, gaps sometimes open in the disk as well. The final outcome of fragment migration can be either tidal disruption or merger into the BH, depending on the fragment radius.
@@
    A major uncertainty in fragment evolution, beyond the scope of this paper to address, concerns the way in which gravitationally unstable fragments will undergo a contraction-collapse sequence to achieve hydrostatic equilibrium. As a limiting case, we consider a situation in which a ($M_{\rm f}\gtrsim 0.1M_{\odot}$) fragment is able to contract to the cold NS mass-radius relationship. This scenario sets a lower limit on the fragment tidal disruption radius and therefore an upper limit on (a) $\gamma$-ray variability from accretion of a disrupted fragment and (b) GW luminosity from fragment inspiral. Under this (optimistic) assumption, we find typical contributions to GRB variability, from the fragment tidal disruption, at the $\sim 10\%$ level, and peak GW strains of $h_{\rm GW}\sim 1.5\cdot 10^{-22}(d/100\penalty 10000\ {\rm Mpc})^{-1}$.
@@ Acknowledgements
Data Availability
The data and code used to make the figures in this paper will be shared upon reasonable request. Inquiries should be directed to Yonatan Lerner (yonatan.lerner@mail.huji.ac.il).
@@ Appendix A Clarification to the CB07 model
[…]
@@ Appendix D Numerical method
$\begin{split}F^{+}-F^{-}=&u^{r}\left[\frac{U_{\rm i}H_{\rm i}-U_{\rm i-1}H_{\rm i-1}}{r_{\rm i}-r_{\rm i-1}}-\right.\vskip 5.69054pt\\ &\qquad-\frac{(U+P)}{\rho}\frac{\left(\rho_{\rm i}H_{\rm i}-\rho_{\rm i-1}H_{\rm i-1}\right)}{r_{\rm i}-r_{\rm i-1}}\biggr]\end{split}$ (45)
@@
Start by choosing: $\alpha$, $a$, $M$, $\dot{M}$, $r_{\rm out}$, an acceptable error ($N_{\rm er}$) and a numerical push ($N_{\rm p}$). Build a logarithmic grid of $N+1$ radial cells, where $r_{\rm i}$ ranges from $r_{\rm ISCO}=r_{\rm N}$ to $r_{\rm out}=r_{\rm 0}$ . Solve for $\eta_{\rm e}$ and $\theta$ at $r_{\rm out}$ using the pressure equation, $Y_{\rm e}=0.5$ and $U/\rho=GM/r_{\rm out}$. Try solving for $\eta_{\rm e}(r_{\rm i}),$ $\theta(r_{\rm i})$, and $Y_{e}(r_{{\rm i}})$ at $r_{{\rm i}}$ with the $\eta_{\rm e}(r_{\rm i-1}),$ $\theta(r_{\rm i-1})$, and $Y_{\rm e}(r_{{\rm i-1}})$ from the previous $r_{{\rm i-1}}>r_{{\rm i}}$ using 3D root-finding, when solving for the pressure, energy and lepton conservation equations. Check whether root-finding worked by testing that for the lepton equation $\Big|\frac{RHS({\rm i})-LHS({\rm i})}{LHS({\rm i})}\Big|<N_{{\rm er}}$ Continue to the next step where $\rm i=i+1$}. Set $Y_{\rm e}(r_{\rm i})=Y_{\rm e}(r_{\rm i-1})+N_{\rm p}$ and solve for $\eta_{\rm e}(r_{\rm i}),$ $\theta(r_{\rm i})$ using a 2D root-finding solving for the pressure and energy equations. Then use the $Y_{\rm e}(r_{\rm i})$ and the $\eta_{\rm e}(r_{\rm i}),$ $\theta(r_{\rm i})$ found in the 2D root-finding to solve for the next step. Where $\rm i=i+1$. Collect all the calculated $\eta_{\rm e}(r_{\rm i}),$ $\theta(r_{\rm i})$, and $Y_{e}(r_{{\rm i}})$ for all $r_{\rm i}$’s. Now we can calculate all of the disks properties, like temperature, density, scale height, pressure and more. While ${\rm i}\leq{\rm N}$   TrueWhen $\rm i>N$False
Figure 15: A flow chart of the numerical algorithm used to solve our model. RHS(i) and LHS(i) are the right and left-hand sides (respectively) of the lepton equation with values from the 3D root finding at step $r_{\rm i}$. $N_{\rm p}$ is a small numerical push set to $10^{-15}$, which was needed because root finding had trouble at the outer regions where changes in $Y_{\rm e}$ are small. $N_{\rm er}=10^{-6}$.
@@ Appendix E Migration calculations
$f_{\rm\gamma}$ $=\begin{cases}1&\tau_{{\rm\nu}}+\tau_{{\rm\widetilde{\nu}}}<1\\ \frac{\sqrt{x/2}+1/\gamma}{\sqrt{x/2}+1}&\tau_{{\rm\nu}}+\tau_{{\rm\widetilde{\nu}}}\geq 1\end{cases}$ (47c)
@@ F.1 GRB variability from tidal disruption of fragment
Figure 16: The ratio between the luminosity from the tidal disruption of the fragment to the steady state luminosity from BH accretion ($L_{\rm TD}/L_{\bullet}$). Left and right panels represent the same $\{M,\dot{M},\alpha\}$ parameter choices as in Fig. 2. White, black, gray and magenta squares are colored with the same color scheme as $1$. Squares that: (1) have a number, represent the ratio between the luminosities as colored by the rainbow color system represented by the bar legend below the figure; (2) are filled in cyan represent that the tidal radius is smaller than the ISCO, meaning no TDE; (3) are filled in light grey, represent final fragment masses below the minimal stable mass for the cold EOS. The luminosity from the tidal disruption is roughly $\sim 10\%$ of the steady luminosity.
@@ F.5 Duration of $\dot{m}_{\rm f}>\dot{M}$
The fragment accretion rate $\dot{m}_{\rm f}$ commonly exceeds the BH accretion rate $\dot{M}$ for a short duration of time (as is visible in the main text, in Figs. 9 and 10). In this appendix we plot (in Fig. 28) two cross section of our parameter survey, in order to explore the duration of this period (which may be observable as extreme GRB variability, as a secondary jet from the accreting fragment may outshine the primary jet from the central BH). Fig. 28shows that this duration generally ranges between 0.511.2ms0.5-11.2ms.
@@ F.6 Extended parameter survey
Figure 29: An extended version of Fig. 6 in the main text, plotting the color-coded values for the largest initial fragment mass (represented as $\log_{10}(m_{\rm f,i}/M_{\odot})$) in the fragmenting $Q<1$ region, but now exploring the full 3 dimensional parameter space of $\{M,\dot{M},\alpha\}$. Each of the panels represents a different $\alpha$ value, ranging from 0.010.30.01-0.3, while varying $\{M,\dot{M}\}$. Black, white, and grey squares represent breakdowns in model assumptions or solutions as in Fig. 4. We see that fragment initial masses span a wide range, from $4\times 10^{-4}\lesssim m_{\rm f}/M_{\odot}\lesssim 2\times 10^{-1}$. The initial masses tend to increase with declining accretion rate, or with increasing central mass or viscosity.
@@
[…]